Determine whether each of the function from 𝑍 to Z is one to one


(a) 𝑓(𝑛) = 𝑛 βˆ’ 1


(b) 𝑓(𝑛) = 𝑛² + 1

1
Expert's answer
2022-03-20T06:41:33-0400

(a) For any "m\in\mathbb{Z}" the equation "f(n)=n-1=m" has a unique solution: "n=m+1". This means that the function "f(n)" is surjective and injective (one-to-one), and it's inverse is "f^{-1}(m)=m+1" .


(b) For any "n\in\mathbb{Z}" "f(n)=f(-n)=n^2+1". Thus, different values of argument may give the same result. Therefore, the function "f(n)" is not one-to-one.


Answer. (a) yes; (b) no.